A certain computer takes an average of 6 seconds to solve a 4- letter matching puzzle. Research has shown that the average time to solve a matching puzzle is directly proportional to the total number of possible choices for the letters. Which of the following is closest to the average length of time it would take to solve an 8-letter matching puzzle? "Letter" refers to the 26 uppercase letters of the English alphabet. Letters may be repeated. Assume the computer works continuously and without interruption.
a) 1 hour
b) 1 day
c) 1 month
d) 1 year

Answer by Boreal(15235)

Possible choices for 4 letters is 26^4=456,976
for 8, it is 26^8=2 x 10^11
That ratio is 456976
time=k*combinations
6=k*456,976
k=6/456976=0.00001313
time for 24= 0.00001313* 2 x 10^11=2,741,899
That is closest to a month
30 days*86,40-0 seconds per day=2,592,000
C